OpenStudy (anonymous):

covert the radian into degrees -405

Nnesha (nnesha):

\[\huge\rm r \times \frac{ 180}{ \pi}\]

OpenStudy (anonymous):

3.14?

Nnesha (nnesha):

that's the formula to convert radians to degrees

OpenStudy (anonymous):

is that answer

Nnesha (nnesha):

3.14 ?? nopwe...

OpenStudy (anonymous):

put r = -405 in the above formula